Brad is 6 years older than Pedro. Five years ago Pedro's age was three-fourths of Brad's age at that time. Find the present ages of Brad and Pedro.
Answer by algebrahouse.com(1659)

x = Pedro's age now
x + 6 = Brad's age now

x - 5 = Pedro's age five years ago
x + 1 = Brad's age five years ago

x - 5 = (3/4)(x + 1) {five years ago Pedro's age was 3/4 Brad's age}
4[x - 5 = (3/4)(x + 1)] {multiplied entire equation by 4 to clear out fractions}
4(x) + 4(-5) = 4(3/4)(x + 1) {multiplied each term by 4}
4x - 20 = 3(x + 1) {multiplied through}
4x - 20 = 3x + 3 {used distributive property}
x - 20 = 3 {subtracted 3x from both sides}
x = 23 {added 20 to both sides}
x + 6 = 29 {substituted 23, in for x, into x + 6}

Pedro is 23 and Brad is 29
www.algebrahouse.com